# sample.bib

### v3

@@ Abstract
Conclusion: These recently released LLMs demonstrated expert-level performance in answering radiation oncology physics questions, exhibiting great potential to assist in radiation oncology physics education and training.
@@ 1 Introduction
Large language models (LLMs) have advanced rapidly. On the one hand, the size of the data used for pre-training and the number of model parameters have grown significantly [gururangan-etal-2020-dont, kaplan2020scalinglawsneurallanguage]; on the other hand, the fine-tuning methods and prompt engineering strategies have advanced substantially [ouyang2022traininglanguagemodelsfollow, hu2021loralowrankadaptationlarge]. Furthermore, agents and Retrieval-Augmented Generation (RAG) systems built on LLMs have seen considerable progress [xi2023risepotentiallargelanguage, lewis2021retrievalaugmentedgenerationknowledgeintensivenlp]. The notable recent developments include OpenAI o1-preview [o1_webpage], GPT-4o [openai2024gpt4ocard], LLaMA 3.1 (405B parameters) [grattafiori2024llama3herdmodels], Gemini 1.5 Pro [geminiteam2024gemini15unlockingmultimodal], and Claude 3.5 Sonnet [claude_webpage], each demonstrating remarkable improvements in language comprehension, reasoning, and domain-specific expertise.
@@
In healthcare, LLMs have been explored for numerous potential applications [Liu2024a, LIU2023100045, zhangkai2024, HOLMES2024e515, li2024, liu2024automatedradiotherapytreatmentplanning]. For their direct use in radiation oncology, unique challenges related to evaluation and validation are posed – due to the high level of complexity and stringent precision required [dawson2007advances, xing2006overview]. This is particularly true for advanced radiation therapy modalities like intensity-modulated X-ray-based radiation therapy and intensity-modulated proton therapy [quan2013preliminary, cao2012uncertainty, zaghian2017comparison, liu2018impact, an2017robust, liu2020robust, shan2020intensity]. Therefore, assessing the performance of LLMs in addressing questions related to radiation oncology physics is crucial. Such evaluations not only tell us how efficiently they process and reason about radiation oncology physics but also help us understand their limitations. This provides us with guidance in this field, especially for their potential utilization in education and training. Built on our prior work [10.3389/fonc.2023.1219326], we provided an updated study with refined methods, offering insights into the performance of the recently released LLMs in radiation oncology physics.
@@ 2 Methods
All the evaluated LLMs were queried with the exam questions through Application Programming Interface (API) services provided by their respective hosts, except LLaMA 3.1 (405B), an open-source LLM, which was hosted by us locally at our institution. All the LLMs used were the recently released version before September 30, 2024. The temperature was set to 0.1 for all LLMs to minimize variability in their responses [^1] The temperature was set to 0.1 rather than 0 due to the different meanings of a temperature of 0 across different LLMs., with the exception of the OpenAI o1-preview, whose temperature was fixed at 1 and could not be changed by the user.
@@ 2.1 Randomly shuffling the answer options
This allowed the LLMs to reason and answer freely. Table 1 illustrates an example of the trials and how we queried the LLMs with the questions. For the responses generated by the LLMs, we utilized the LLaMA 3.1 (405B) model hosted locally to further extract the chosen answer options (letters A, B, C, D, or E) from responses. We then conducted manual verification and correction of the extracted options and obtained the final answer option sheet for all LLMs. The accuracy of each LLM was reported as the mean score across the five trials, and the measurement uncertainty was reported as the standard deviation of the five trials, as shown in the following equations:
@@
Trail Prompt Question
Trail 1 Please solve this radiation oncology physics problem:
@@ 2.2.1 Replacing the correct option with ”None of the above”
As before, we utilized the previously described processes for answer-option extraction, manual verification, and correction, as outlined in Sec. 2.1. The performance accuracy and uncertainty of each LLM were reported as the average score and standard deviation across all five trials. Due to this setup, these exams were not used to test humans, as this pattern can be easily recognized by human test-takers [^2] Since each question was queried through the API individually, it is assumed that LLMs would not notice this pattern.
@@ 2.2.2 Explain-first and step-by-step instruction
To further check if explicitly asking the LLMs to explain first and then develop answers step-by-step would improve their deductive reasoning ability [wei2023chainofthoughtpromptingelicitsreasoning], we engineered the following prompt and queried the LLMs again with it:
@@ 3.1 Results of exam sets with randomly shuffled options
The raw counts of incorrect responses by the LLMs are shown in Fig. 2, where each color represents the incorrect answers by an LLM across trials. As observed, each LLM exhibited variability in answering questions across trials. Notably, the models also showed similarities in incorrectly answering certain questions. We analyzed the questions that all five models had ever answered incorrectly across all five trials – question numbers: 14, 27, 33, 42, 67, 87, 95, 96. Interestingly, only one of these questions was math-based, while the remaining seven were closely related to clinical medical physics knowledge, such as American Association of Physicists in Medicine (AAPM) Task Group (TG) reports and clinical experience. This observation suggests that current LLMs may still struggle with answering clinically focused radiation oncology physics questions. For example, question number 42 does not involve any calculations but instead focuses primarily on clinical hands-on experience.
Figure 2: Distribution of incorrect answer counts of each LLM across five exam sets where answer options were randomly shuffled. Questions that were correctly answered by LLMs in all five trials are not displayed in this figure.
@@ 3.2 Results of LLMs’ deductive reasoning ability
Fig. 3 shows the results of the deductive reasoning ability tests, where the correct answer options were replaced with ”None of the above” in all questions. Overall, all LLMs performed much more poorly compared to the results in Sec. 3.1. Given that transformer-based LLMs [vaswani2023attentionneed] were designed to predict the next word in a sequence, replacing the correct answers with ”None of the above” would likely disrupt their pattern recognition abilities, thereby reducing their overall scores performed on the exam sets. Nonetheless, the OpenAI o1-preview and GPT-4o still outperformed the others, especially on math-based questions, indicating the strong reasoning ability of these two models.
@@
Fig. 4 compares the performance of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet models with the original simple prompts and with the explain-first, step-by-step instruction prompts. Overall, all three models demonstrated improved reasoning ability with the latter prompting strategy. Notably, Gemini 1.5 Pro showed significant gains on math-based questions, increasing its score from 24% to 68%. The o1-preview and GPT-4o improved by only about 1% overall, which is too small to be represented in this figure.
Figure 4: Improvements in reasoning abilities observed using the explain-first and step-by-step instruction prompts for the LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. Figure (a) illustrates the improvements for all questions and math-based questions, while Figure (b) presents the improvements broken down by different topics. (Note: The o1-preview and GPT-4o showed only about 1% overall improvement, too small to be represented in this figure.)
@@ 4 Discussion
Over the past two years, our studies have observed a notable improvement in the performance of state-of-the-art LLMs on this highly specialized task – answering radiation oncology physics questions, as shown in Fig. 5. Early versions of ChatGPT, like GPT-3.5 in late 2022 [chatgpt_webpage], scored around 54%, showing clear gaps in domain-specific knowledge. With the introduction of GPT-4 in early 2023 [openai2024gpt4technicalreport], performance jumped to around 76%, reflecting improvements in accuracy and understanding. Subsequent releases of the GPT-4o model and more recently the o1-preview (both in 2024), pushed scores even higher to 90% and 94% respectively, indicating increasing capabilities in radiation oncology physics. This steady improvement can be attributed to more extensive domain pre-training, refined architectural updates [brown2020languagemodelsfewshotlearners], and enhanced fine-tuning techniques, all of which have led to stronger reasoning skills, improved understanding, and better alignment with expert-level knowledge. The evolution of these models over the last two years underscores the rapid growth of LLMs, suggesting their potential as useful tools in areas such as radiation oncology physics education and training.
@@
4.2 Utilization of LLMs in radiation oncology physics training and education
With the excellent performance of LLMs in radiation oncology physics evaluated in this study, it offers a promising avenue for their utilization in radiation oncology physics education to enhance learning efficiency and accessibility. For example, LLMs can assist in educating graduate students, postdocs, and medical physics residents as they prepare for challenging medical physics certification exams, such as the American Board of Radiology certification exams. LLMs have the unique advantage of providing immediate and context-specific explanations of medical physics concepts and are efficient at retrieving information from a wide range of topics. In busy clinical environments, where mentors may sometimes be occupied with clinical duties, LLMs can serve as a useful tool to address easy questions and have trainees pay more attention to more complex problems. This leads to more focused and productive discussions with mentors. In summary, LLMs can bridge gaps in traditional learning settings and support self-directed study in radiation oncology physics.
@@ 5 Conclusion
We evaluated recently released LLMs using a method that randomly shuffled the answer options of radiation oncology physics questions. Our results demonstrated that these models achieved expert-level performance on these questions, with some even surpassing human experts with a majority vote. However, when the correct answer options were replaced with ”None of the above,” all models exhibited a steep decline in performance, suggesting room for improvement. Employing the technique of explain-first and step-by-step instruction enhanced the reasoning abilities of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. The performance of these recently released LLMs exhibited great potential for assisting in radiation oncology physics education and training.

### v4

@@ Abstract
Conclusion: These recently released LLMs demonstrated expert-level performance in answering radiation oncology physics questions.
@@ 1 Introduction
Large language models (LLMs) have advanced rapidly. On the one hand, the size of the data used for pre-training and the number of model parameters have grown a lot. For example, GPT-2 had 1.5 billion parameters [gpt2_webpage], GPT-3 scaled up to 175 billion [NEURIPS2020_1457c0d6], and GPT-4 is estimated to have even more [openai2024gpt4technicalreport]. On the other hand, the fine-tuning methods and prompt engineering strategies have advanced substantially [NEURIPS2022_b1efde53, hu2021loralowrankadaptationlarge]. Furthermore, agents and Retrieval-Augmented Generation (RAG) systems built on LLMs have seen considerable progress [scichinaiformatic, NEURIPS2020_6b493230]. Notable recent developments as of September 2024 include OpenAI o1-preview [o1_webpage], GPT-4o [openai2024gpt4ocard], LLaMA 3.1 (405B parameters) [grattafiori2024llama3herdmodels], Gemini 1.5 Pro [geminiteam2024gemini15unlockingmultimodal], and Claude 3.5 Sonnet [claude_webpage], demonstrating state-of-art performance in overall language processing, reasoning, and diverse downstream applications.
@@
In healthcare, LLMs have been explored for numerous potential applications [Liu2024a, LIU2023100045, zhangkai2024, HOLMES2024e515, li2024]. For their direct use in radiation oncology, unique challenges related to evaluation and validation arise due to the complexity and precision of treatment, which involves both clinical factors and physics considerations. Therefore, assessing the performance of LLMs in addressing questions related to radiation oncology physics is crucial. Such evaluations not only tell us how efficiently they process and reason about radiation oncology physics but also help us understand their limitations. In the past, several LLMs were evaluated on the 2021 American College of Radiology (ACR) Radiation Oncology In-Training Examination (TXIT), revealing that GPT-4-turbo achieved the highest score of 68.0%, outperforming some resident physicians [doi:10.1089/aipo.2023.0007]. GPT-3.5 and GPT-4 were also assessed on Japan’s medical physicist board examinations from 2018 to 2022, where GPT-4 demonstrated superior performance with an average accuracy of 72.7% [Kadoya2024-pi]. To offer insights into the recently released state-of-art LLMs and build on our prior work [10.3389/fonc.2023.1219326], we present here an updated study with refined methods evaluating their performance in radiation oncology physics.
@@ 2 Methods
All the evaluated LLMs were queried with the exam questions through Application Programming Interface (API) services provided by their respective hosts, except LLaMA 3.1 (405B), an open-source LLM, which was hosted by us locally at our institution. All the LLMs used were the recently released version before September 30, 2024. The temperature was set to 0.1 for all LLMs to minimize variability in their responses [^1] The temperature was set to 0.1 rather than 0 due to the different meanings of a temperature of 0 across different LLMs. To avoid potential unexpected behaviors from models, we set the lower bound to 0.1 rather than 0. , with the exception of the OpenAI o1-preview, whose temperature was fixed at 1 and could not be changed by the user.
@@ 2.1 Randomly shuffling the answer options
This allowed the LLMs to reason and answer freely. Table 1 illustrates an example of the trials and how we queried the LLMs with the questions. For the responses generated by the LLMs, we utilized the LLaMA 3.1 (405B) model hosted locally to further extract the chosen answer options (letters A, B, C, D, or E) from free-form responses, thereby reducing some of the manual effort required to read and record them individually. We then conducted manual verification of the extracted options and obtained the final answer option sheet for all LLMs to compare with the ground truth answers. The accuracy of each LLM was reported as the mean score across the five trials, and the measurement uncertainty was reported as the standard deviation of the five trials, as shown in the following equations:
@@
Trial Prompt Question
Trial 1 Please solve this radiation oncology physics problem:
@@ 2.2.1 Replacing the correct option with ”None of the above”
As before, we utilized the previously described processes for answer-option extraction and manual verification, as outlined in Sec. 2.1. The performance accuracy and uncertainty of each LLM were reported as the average score and standard deviation across all five trials. Due to this setup, these exams were not used to test humans, as this pattern can be easily recognized by human test-takers [^2] Since each question was queried through the API individually, it is assumed that LLMs would not notice this pattern..
@@ 2.2.2 Explain-first and step-by-step instruction
To further check if explicitly asking the LLMs to explain first and then develop answers step-by-step (chain-of-thought) would improve their deductive reasoning ability [wei2023chainofthoughtpromptingelicitsreasoning], we engineered the following prompt and queried the LLMs again with it:
@@ 3.1 Results of exam sets with randomly shuffled options
The raw counts of incorrect responses by the LLMs are shown in Fig. 2, where each color represents the incorrect answers by an LLM across trials. As observed, each LLM exhibited variability in answering questions across trials. Notably, the models also showed similarities in incorrectly answering certain questions. We analyzed the questions that were commonly answered incorrectly by all LLMs at least once across all five trials – question numbers: 14, 27, 42, 67, 87, 95, and 96. Interestingly, only one of these questions was math-based, while the remaining seven were closely related to clinical medical physics knowledge, such as American Association of Physicists in Medicine (AAPM) Task Group (TG) reports and clinical experience. This observation suggests that current LLMs may still struggle with answering clinically focused radiation oncology physics questions. For example, question number 42 does not involve any calculations but instead focuses primarily on clinical hands-on experience.
Figure 2: Distribution of incorrectly selected answer options by each LLM across five exam sets. All answer options in the five exam sets were randomly shuffled. Questions that were correctly answered by all LLMs in all five trials are not shown in this figure. Questions 14, 27, 42, 67, 87, 95, and 96, which were commonly answered incorrectly by all LLMs at least once, are underlined in the figure.
@@ 3.2 Results of LLMs’ deductive reasoning ability
Fig. 3 shows the results of the deductive reasoning ability tests, where the correct answer options were replaced with ”None of the above” in all questions. Overall, all LLMs performed much more poorly compared to the results in Sec. 3.1. Given that transformer-based LLMs [NIPS2017_3f5ee243] were designed to predict the next word in a sequence, replacing the correct answers with ”None of the above” would likely disrupt their pattern recognition abilities, thereby reducing their overall scores performed on the exam sets. Nonetheless, the OpenAI o1-preview and GPT-4o still outperformed the others, especially on math-based questions, indicating the strong reasoning ability of these two models.
Fig. 4 compares the performance of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet models with the original simple prompts and with the explain-first, step-by-step instruction prompts. Overall, all three models demonstrated improved reasoning ability with the latter prompting strategy. Notably, Gemini 1.5 Pro showed significant gains on math-based questions, increasing its score from 24% to 68%. The o1-preview and GPT-4o showed only about a 1% overall difference, which was too small to be represented in this figure.
@@
Figure 4: Comparison of accuracy between the short prompt and the explain-first, step-by-step instruction prompt (chain-of-thought) for LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet. Figure (a) shows the comparison between the short prompt and the explain-first, step-by-step instruction prompt for all questions and math-based questions, while Figure (b) breaks down the comparison by topic. (Note: o1-preview and GPT-4o showed only about a 1% overall difference, which was too small to be represented in this figure.)
@@ 4 Discussion
Over the past two years, our studies have observed a notable improvement in the performance of state-of-the-art LLMs on this highly specialized task – answering radiation oncology physics questions, as shown in Fig. 5. Early versions of ChatGPT, like GPT-3.5 in late 2022 [chatgpt_webpage], scored around 54%, showing clear gaps in domain-specific knowledge. With the introduction of GPT-4 in early 2023, performance jumped to around 76%, reflecting improvements in accuracy and understanding. Subsequent releases of the GPT-4o model and more recently the o1-preview (both in 2024), pushed scores even higher to 90% and 94% respectively, indicating increasing capabilities in radiation oncology physics. This steady improvement can be attributed to more extensive domain pre-training, increase of number of parameters, refined architectural updates, and enhanced fine-tuning techniques [gururangan-etal-2020-dont, kaplan2020scalinglawsneurallanguage], all of which have led to improved understanding, stronger reasoning skills, and better alignment with expert-level knowledge. The evolution of these models over the last two years underscores the rapid growth of LLMs, suggesting their potential as useful tools in areas such as radiation oncology physics education and training.
@@
4.2 Potential applications of LLMs in radiation oncology physics
Recent advancements in exploring potential applications of LLMs in radiation oncology physics have focused on auto-contouring, dose prediction and treatment planning. For auto-contouring, LLMs have been utilized to extract electronic medical records (EMR) text data and align them with the image embeddings of the mixture-of-experts model to improve the performance of the target volume contouring for radiation therapy [RAJENDRAN2025230]. In addition, LLMs have also been used to extract text-based features and incorporated them into vision transformer to help improve the target delineation results [oh2024mixturemulticenterexpertsmultimodal]. In dose prediction, LLM have been explored to encode knowledge from prescriptions and interactive instructions from clinicians into neural networks to enhace the prediction of dose-volume histograms (DVH) from medical images [https://doi.org/10.1002/mp.17416]. Regarding treatment planning, GPT-4V has been investigated for evaluating dose distribution and DVH and assisting with the optimization of the treatment planning [liu2025automatedradiotherapytreatmentplanning]. Furthermore, an LLM-based multi-agent system has also been developed to mimick the workflow of dosimetrists and medical physicists to generate text-based treatment plans [Wang_2025]. Collectively, these advancements highlight the transformative potential of LLMs in radiation oncology physics, offering potential improvements in efficiency and outcomes.
4.3 Possible further improvement of LLMs in radiation oncology physics
The performance of LLMs on radiation oncology physics, although encouraging, still requires further improvement due to two primary factors. First, radiation oncology physics represents a very specialized domain characterized by both the complexity of physics concepts and specific clinical contexts, neither of which was extensively represented in the general datasets used during the initial pre-training of these models. Second, existing LLMs still encounter difficulties with reasoning tasks specific to radiation oncology physics, indicating a need for enhanced general reasoning capabilities. To address these limitations, further studies could explore strategies of fine-tuning existing LLMs using specialized medical physics domain datasets with clinical contexts. Such fine-tuning would likely enable the models to better capture the complexities and contextual details of the domain, enhancing their accuracy and practical clinical utility in medical physics tasks. Additionally, to improve reasoning capabilities, techniques such as chain-of-thought, which encourages models to articulate intermediate reasoning steps explicitly, and reinforcement learning, which optimizes model responses in desired patterns, could be investigated [shao2024deepseekmathpushinglimitsmathematical].
4.4 Limitations
Although the LLMs evaluated in this study exhibit expert-level performance on radiation oncology physics questions, such results do not directly translate to effectiveness in practical clinical tasks like treatment planning and delivery. This limitation arises from differences between theoretical examinations and practical clinical applications. Clinical scenarios encountered in radiation oncology are inherently more complex, context-dependent, and require integrating multiple sources of clinical and patient-specific data, whereas theoretical examinations often have clearly defined questions and objective answers. Consequently, strong performance in controlled question-answering tasks may not effectively transfer to real-world contexts, which frequently involve ambiguity, uncertainty, and nuanced clinical judgment. Additionally, clinical decision-making encompasses not only physics-based calculations but also multidisciplinary collaboration, patient safety considerations, regulatory compliance, and human factors in clinical workflows. Therefore, although the evaluated models demonstrate promise in foundational physics knowledge, caution must be exercised when inferring their direct clinical utility.
@@ 5 Conclusion
We evaluated recently released LLMs using a method that randomly shuffled the answer options of radiation oncology physics questions. Our results demonstrated that these models achieved expert-level performance on these questions, with some even surpassing human experts with a majority vote. However, when the correct answer options were replaced with ”None of the above,” all models exhibited a steep decline in performance, suggesting room for improvement. Employing the technique of explain-first and step-by-step instruction prompt enhanced the reasoning abilities of LLaMA 3.1 (405B), Gemini 1.5 Pro, and Claude 3.5 Sonnet.